OpenStudy (anonymous):

If a, x, and b are consecutive terms of a geometric sequence, then x is called the geometric mean of a and b. Find the geometric mean of 2 and 8.

OpenStudy (anonymous):

geometric mean is \[\sqrt{ab}\]

OpenStudy (anonymous):

I thought it had the formula tn = ar^n-1

OpenStudy (anonymous):

so in your case it would be \[\sqrt{2\times 8}=\sqrt{16}=4\]

OpenStudy (anonymous):

if you have a geometric sequence \[2,2r, 2r^2,...\] and you know that \[2r^2=8\] that means \[r^2=4\] \[r=2\] and so the middle term \[2r=4\]

OpenStudy (anonymous):

ohhh, i see. thanks!

OpenStudy (anonymous):

yw